### v1

# thanks: Corresponding Author

Abstract
Axial-vector hadronic molecules $\mathcal{M}_{\mathrm{AV}}=\Upsilon B_{c}^{-}$ and $\widetilde{\mathcal{M}}_{\mathrm{AV}}=\eta_{b}B_{c}^{\ast-}$ with the quark content $bb\overline{b}\overline{c}$ are studied using QCD sum rule method. The spectroscopic parameters of these molecules are computed in the context of the two-point sum rule method. Predictions for their masses are identical to each other and confirm that they are structures unstable against dissociations to ordinary heavy mesons. We evaluate the width of the state $\mathcal{M}_{\mathrm{AV}}$ and assume that it is equal to that of $\widetilde{\mathcal{M}}_{\mathrm{AV}}$. To this end, we explore its dominant decay channels $\mathcal{M}_{\mathrm{AV}}\to\Upsilon B_{c}^{-}$ and $\mathcal{M}_{\mathrm{AV}}\to\eta_{b}B_{c}^{\ast-}$. There also are subleading modes of $\mathcal{M}_{\mathrm{AV}}$ generated due to annihilation of $\overline{b}b$ quarks. We consider decays of the molecule $\mathcal{M}_{\mathrm{AV}}$ to pairs of the mesons $B^{\ast-}\overline{D}^{0}$, $\overline{B}^{\ast 0}D^{-}$, $B^{-}\overline{D}^{\ast 0}$, $\overline{B}^{0}D^{\ast-}$, $\overline{B}_{s}^{\ast 0}D_{s}^{-}$, and $\overline{B}_{s}^{0}D_{s}^{\ast-}$. To find strong couplings at the $\mathcal{M}_{\mathrm{AV}}$-meson-meson vertices which determine the partial widths of these processes, we apply QCD three-point sum rule approach. The mass $m=(15800\pm 90)~\mathrm{MeV}$ and width $\Gamma[\mathcal{M}_{\mathrm{AV}}]=(112\pm 17)~\mathrm{MeV}$ of the molecule $\mathcal{M}_{\mathrm{AV}}$ are useful for experimental studies of fully heavy molecular structures at ongoing and planning experiments.
I Introduction
It is known that hadronic molecular structures are among objects interesting for high energy physics. These exotic states are presumably composed of ordinary mesons and establish a class in the hadron spectroscopy. Though existence of different hadronic molecules was already supposed Bander:1975fb ; Voloshin:1976ap ; DeRujula:1976zlg , binding mechanisms of these structures, the processes at which they might be discovered and methods for computations of the parameters of such states were elaborated mainly in the later publications Tornqvist:1991lks ; Ding:2008mp ; Zhang:2009vs ; Sun:2012sy ; Chen:2015ata ; Karliner:2015ina ; Liu:2016kqx ; Chen:2017vai ; Sun:2018zqs ; Albuquerque:2012rq ; PavonValderrama:2019ixb ; Molina:2020hde ; Xu:2020evn ; Xin:2021wcr ; Agaev:2022duz ; Agaev:2023eyk ; Braaten:2023vgs ; Wu:2023rrp ; Liang:2023jxh ; Wang:2025zss ; Braaten:2024tbm .
States that are built of only heavy mesons form an interesting and rapidly growing subclass of the hadronic molecules. They may contain an equal number of charm or bottom quarks-antiquarks. This group of molecules was studied in Refs. Agaev:2023ruu ; Agaev:2023rpj ; Yalikun:2025ssz ; Liu:2024pio . We also explored the fully heavy molecules $\eta_{c}\eta_{c}$, $\chi_{c0}\chi_{c0}$, and $\chi_{c1}\chi_{c1}$ by calculating their masses and decay widths Agaev:2023ruu ; Agaev:2023rpj . In these articles we suggested their interpretation as possible candidates of new four $X$ resonances which supposedly are $cc\overline{c}\overline{c}$ particles LHCb:2020bwg ; Bouhova-Thacker:2022vnt ; CMS:2023owd .
Hidden charm-bottom molecules belong to this group of particles as well, and were investigated in Refs. Liu:2024pio ; Liu:2023gla ; Wang:2023bek ; Agaev:2025wdj ; Agaev:2025fwm ; Agaev:2025nkw . Thus, in the framework of the coupled-channel unitary approach the properties of molecules $B_{c}^{(\ast)+}B_{c}^{(\ast)-}$ were addressed in Ref. Liu:2023gla . The parameters and decay channels of the states $B_{c}^{+}B_{c}^{-}$, $(B_{c}^{\ast+}B_{c}^{-}+B_{c}^{+}B_{c}^{\ast-})/2$ and $B_{c}^{\ast+}B_{c}^{\ast-}$ with spin-parities $J^{\mathrm{PC}}=0^{++}$, $1^{++}$, and $2^{++}$ were explored in our publications Agaev:2025wdj ; Agaev:2025fwm ; Agaev:2025nkw .
Hadronic molecules with the nonsymmetrical contents $bb\overline{b}\overline{c}$ and $cc\overline{c}\overline{b}$ establish another group of fully heavy particles. The molecules with these contents also attracted interest of researches Liu:2024pio ; Agaev:2025did . In fact, in Ref. Agaev:2025did we calculated the mass and decay width of the scalar states $\mathcal{M}_{\mathrm{b}}=\eta_{b}B_{c}^{-}$ and $\mathcal{M}_{\mathrm{c}}=\eta_{c}B_{c}^{+}$: Our results demonstrated that they are relatively broad compounds.
In this article, we investigate the axial-vector hadronic molecules $\mathcal{M}_{\mathrm{AV}}=\Upsilon B_{c}^{-}$ and $\widetilde{\mathcal{M}}_{\mathrm{AV}}=\eta_{b}B_{c}^{\ast-}$ and calculate their masses and full decay widths. These structures composed of $bb\overline{b}\overline{c}$ quarks are counterparts of the scalar molecule $\mathcal{M}_{\mathrm{b}}$ from our previous paper Agaev:2025did . Here, computations are performed in the context of QCD SR method Shifman:1978bx ; Shifman:1978by . Parameters of $\mathcal{M}_{\mathrm{AV}}$ and $\widetilde{\mathcal{M}}_{\mathrm{AV}}$ show that despite different internal organizations, they have almost identical masses: There is a difference around of a few $\mathrm{MeV}$ which is considerably smaller than the accuracy of the sum rule method. Therefore, we treat the molecules $\mathcal{M}_{\mathrm{AV}}$ and $\widetilde{\mathcal{M}}_{\mathrm{AV}}$ as identical particles, and concentrate on properties of $\mathcal{M}_{\mathrm{AV}}$.
The structure $\mathcal{M}_{\mathrm{AV}}$ is strong-interaction unstable particle and easily dissociates to constituent mesons $\Upsilon B_{c}^{-}$. The decay to a pair $\eta_{b}B_{c}^{\ast-}$ is also among its kinematically allowed decay channels. These two processes are dominant decay modes of $\mathcal{M}_{\mathrm{AV}}$, in which all quarks from $\mathcal{M}_{\mathrm{AV}}$ participate in formation of the final-state mesons with required spin-parities. Besides, there are channels which are triggered by annihilation of $b\overline{b}$ quarks and generation of light quark-antiquark pairs $\overline{q}q$ and $\overline{s}s$. Having combined with the remaining heavy quarks they create conventional meson pairs. In this work, we explore production of mesons $B^{\ast-}\overline{D}^{0}$, $\overline{B}^{\ast 0}D^{-}$, $B^{-}\overline{D}^{\ast 0}$$\overline{B}^{0}D^{\ast-}$, $\overline{B}_{s}^{\ast 0}D_{s}^{-}$, and $\overline{B}_{s}^{0}D_{s}^{\ast-}$ due to this mechanism. This mechanism for transformation of four-quark states was considered in Refs. Becchi:2020mjz ; Becchi:2020uvq ; Agaev:2023ara and applied to decays of various diquark-antidiquark states. It was successfully used to study processes with molecular structures as well Agaev:2025wdj ; Agaev:2025fwm ; Agaev:2025nkw ; Agaev:2025did .
Partial widths all of these channels depend on numerous parameters such as masses and current couplings (decay constants) of particles involved into a process. They contain also strong couplings of particles at the relevant molecule-meson-meson vertices. These couplings are objects of our present calculations. They can be estimated by applying technical tools of QCD three-point SR method. But there are differences in treatments of the dominant and subleading modes. Thus, in the case of the dominant channels, we apply the standard techniques to compute a correlation function of interest, whereas the subleading decays require additionally a replacement the vacuum expectation value $\langle\overline{b}b\rangle$ by the gluon condensate $\langle\alpha_{s}G^{2}/\pi\rangle$. Nevertheless this operation does not imply introduction into analysis of new parameters.
This paper is structured in the following from: In Sec. II, we calculate the spectroscopic parameters of the axial-vector molecules $\mathcal{M}_{\mathrm{AV}}$ and $\widetilde{\mathcal{M}}_{\mathrm{AV}}$. The Sec. III is devoted to investigation of the dominant decay channels $\mathcal{M}_{\mathrm{AV}}\rightarrow\Upsilon B_{c}^{-}$ and $\mathcal{M}_{\mathrm{AV}}\rightarrow\eta_{b}B_{c}^{\ast-}$. The subleading decays of the molecule $\mathcal{M}_{\mathrm{AV}}$ are considered in Sec. IV. Here, we also find the full decay width of $\mathcal{M}_{\mathrm{AV}}$. The last section V contains our analysis of the results and a few final notes.
II Spectroscopic parameters of the molecules $\mathcal{M}_{\mathrm{AV}}$ and $\widetilde{\mathcal{M}}_{\mathrm{AV}}$
To evaluate the masses and current couplings (pole residues) of the molecular structures $\mathcal{M}_{\mathrm{AV}}$ and $\widetilde{\mathcal{M}}_{\mathrm{AV}}$, we introduce the relevant interpolating currents
$J_{\mu}(x)=\overline{b}_{a}(x)\gamma_{\mu}b_{a}(x)\overline{c}_{b}(x)i\gamma_{5}b_{b}(x),$ (1)
and
$\widetilde{J}_{\mu}(x)=\overline{b}_{a}(x)i\gamma_{5}b_{a}(x)\overline{c}_{b}(x)\gamma_{\mu}b_{b}(x).$ (2)
Here, $b(x)$ and $c(x)$ are the quark fields, whereas $a$ and $b$ denote the color indices.
The SRs for the mass $m$ and current coupling $\Lambda$ of $\mathcal{M}_{\mathrm{AV}}$ can be extracted from analysis of the following correlation function
$\Pi_{\mu\nu}(p)=i\int d^{4}xe^{ipx}\langle 0|\mathcal{T}\{J_{\mu}(x)J_{\nu}^{{\dagger}}(0)\}|0\rangle,$ (3)
with $\mathcal{T}$ being the time-ordered product of two currents.
In the framework of the sum rule method the correlator $\Pi_{\mu\nu}(p)$ should be calculated by two alternative ways. Thus, it has to be found using the parameters $m$ and $\Lambda$. The correlation function calculated by this manner forms the physical side $\Pi_{\mu\nu}^{\mathrm{Phys}}(p)$ of the corresponding SRs. The correlator $\Pi_{\mu\nu}^{\mathrm{Phys}}(p)$ is determined by the expression
$\Pi_{\mu\nu}^{\mathrm{Phys}}(p)$ $=$ $\frac{\langle 0|J_{\mu}|\mathcal{M}_{\mathrm{AV}}(p,\epsilon)\rangle\langle\mathcal{M}_{\mathrm{AV}}(p,\epsilon)|J_{\nu}^{\dagger}|0\rangle}{m^{2}-p^{2}}$ (4)
$+\cdots,$
where $\epsilon_{\mu}$ is the polarization vector of the axial-vector molecule. The term presented above explicitly is the contribution of the ground-level molecule, whereas ellipses denote contributions of the higher resonances and continuum states.
We calculate $\Pi_{\mu\nu}^{\mathrm{Phys}}(p)$ by introducing the matrix element
$\langle 0|J_{\mu}|\mathcal{M}_{\mathrm{AV}}(p,\epsilon)\rangle=\Lambda\epsilon_{\mathcal{\mu}}(p),$ (5)
and obtain
$\Pi_{\mu\nu}^{\mathrm{Phys}}(p)=\frac{\Lambda^{2}}{m^{2}-p^{2}}\left(g_{\mu\nu}-\frac{p_{\mu}p_{\nu}}{m^{2}}\right)+\cdots.$ (6)
The correlation function contains two Lorentz structures. Because the term $\sim g_{\mu\nu}$ receives contributions arising only from the spin-1 particle, we choose it to carry our the sum rule analysis. Then the factor $\Lambda^{2}/(m^{2}-p^{2})$ is the invariant amplitude $\Pi^{\mathrm{Phys}}(p^{2})$ necessary for future studies.
The correlator $\Pi_{\mu\nu}(p)$ should be computed by employing heavy quark propagators $S_{b(c)}(x)$ and using techniques of the operator product expansion ($\mathrm{OPE}$) . As a result, we get
[…]
$\left.\times S_{b}^{ba^{\prime}}(x)\gamma_{\nu}S_{b}^{a^{\prime}a}(-x)\right]-\mathrm{Tr}\left[\gamma_{\mu}S_{b}^{aa^{\prime}}(x)\gamma_{\nu}S_{b}^{a^{\prime}a}(-x)\right]$
$\left.\times\mathrm{Tr}\left[\gamma_{5}S_{b}^{bb^{\prime}}(x)\gamma_{5}S_{c}^{b^{\prime}b}(-x)\right]\right\}.$ (7)
The explicit expressions for $S_{b(c)}(x)$ can be found in Ref. Agaev:2020zad . The function $\Pi_{\mu\nu}^{\mathrm{OPE}}(p)$ establishes the QCD component of the sum rules.
The function $\Pi_{\mu\nu}^{\mathrm{OPE}}(p)$ is also composed of two terms. We choose a term which is proportional to $g_{\mu\nu}$ and label by $\Pi^{\mathrm{OPE}}(p^{2})$ corresponding invariant amplitude. Afterwards, we equate amplitudes $\Pi^{\mathrm{Phys}}(p^{2})$ and $\Pi^{\mathrm{OPE}}(p^{2})$, carry out required operations detailed in Ref. Agaev:2024uza , and obtain the sum rules for parameters $m$ and $\Lambda$ of the molecule $\mathcal{M}_{\mathrm{AV}}$
$m^{2}=\frac{\Pi^{\prime}(M^{2},s_{0})}{\Pi(M^{2},s_{0})},$ (8)
and
$\Lambda^{2}=e^{m^{2}/M^{2}}\Pi(M^{2},s_{0}).$ (9)
Above, we use the notation $\Pi^{\prime}(M^{2},s_{0})=d\Pi(M^{2},s_{0})/d(-1/M^{2})$, where $\Pi(M^{2},s_{0})$ is the amplitude $\Pi^{\mathrm{OPE}}(p^{2})$ subjected to Borel transformation and continuum subtraction procedures. The first of them is done to subdue effects of higher resonances and continuum states, whereas the second procedure permits one to subtract these terms from the QCD side of the SR formula. Then, $\Pi(M^{2},s_{0})$ becomes a function of the Borel $M^{2}$ and continuum subtraction $s_{0}$ parameters. In the case under analysis, it is given by the formula
$\Pi(M^{2},s_{0})=\int_{(3m_{b}+m_{c})^{2}}^{s_{0}}ds\rho^{\mathrm{OPE}}(s)e^{-s/M^{2}}+\Pi(M^{2}).$ (10)
Here, $\rho^{\mathrm{OPE}}(s)$ is the spectral density found as an imaginary part of the function $\Pi^{\mathrm{OPE}}(p^{2})$. In the current work we take into account contributions to $\Pi^{\mathrm{OPE}}(p^{2})$ coming from the perturbative and dimension-four $\sim\langle\alpha_{s}G^{2}/\pi\rangle$ terms , therefore $\rho^{\mathrm{OPE}}(s)$ consists of $\rho^{\mathrm{pert.}}(s)$ and $\rho^{\mathrm{Dim4}}(s)$ components. The function $\Pi(M^{2})$ is extracted from the correlation function $\Pi^{\mathrm{OPE}}(p)$ and does not contain terms included into $\rho^{\mathrm{OPE}}(s)$.
To perform computations, one should specify the parameters in Eqs. (8) and (9). The gluon condensate $\langle\alpha_{s}G^{2}/\pi\rangle=(0.012\pm 0.004)~\mathrm{GeV}^{4}$ as well as masses $m_{b}=(4.183\pm 0.007)~\mathrm{GeV}$ and $m_{c}=(1.2730\pm 0.0046)~\mathrm{GeV}$ of the quarks are universal parameters Shifman:1978bx ; Shifman:1978by ; PDG:2024 . Contrary, the pair $M^{2}$ and $s_{0}$ is fixed by a problem under consideration and should meet well-known constraints of SR investigations. One of them is dominance of the pole contribution ($\mathrm{PC}$) to obtained quantities: As a result, one demands fulfilment of the restriction $\mathrm{PC}\geq 0.5$ . Convergence of $\mathrm{OPE}$ is the next constraint of credible SR investigations. Because, the correlator $\Pi(M^{2},s_{0})$ contains dimension-$4$ nonperturbative term $\Pi^{\mathrm{Dim4}}(M^{2},s_{0})$, it is enough fulfillment of $|\Pi^{\mathrm{Dim4}}(M^{2},s_{0})|\leq 0.05|\Pi(M^{2},s_{0})|$ which guarantees convergence of $\mathrm{OPE}$. One has also to remember that extracted quantities should be stable against variations of $M^{2}$ and $s_{0}$.
Armed with this knowledge, we have performed numerical computations of $m$ over wide range of the parameters $M^{2}$ and $s_{0}$: Some of these computations are visualized in Fig. 1.
Figure 1: The mass $m$ as a function of parameters $M^{2}$ at various $s_{0}$. Two vertical lines fix borders of $M^{2}$ inside of which restrictions imposed on $\Pi(M^{2},s_{0})$ are fulfilled.
Collected predictions allow us to find the windows for $M^{2}$ and $s_{0}$, where all restrictions discussed above are satisfied. We find that intervals
$M^{2}\in[16,19]~\mathrm{GeV}^{2},\ s_{0}\in[278,283]~\mathrm{GeV}^{2},$ (11)
meet these constraints. Indeed, at maximal $M^{2}$ the $s_{0}$-averaged pole contribution is $\mathrm{PC}\approx 0.51$, whereas at minimal value of the Borel parameter it is equal to $\mathrm{PC}\approx 0.63$. At $M^{2}=16~\mathrm{GeV}^{2}$ the nonperturbative term is negative and forms approximately $1.4\%$ of $\Pi(M^{2},s_{0})$. As a function of $M^{2}$ the $\mathrm{PC}$ is depicted in Fig. 2, where all lines exceed the limit $\mathrm{PC}=0.5$.
Figure 2: Dependence of $\mathrm{PC}$ on the Borel parameter $M^{2}$ at fixed $s_{0}$. The circle shows the point $M^{2}=17.5~\mathrm{GeV}^{2}$ and $s_{0}=280.5~\mathrm{GeV}^{2}$.
We extract $m$ and $\Lambda$ as mean values of these parameters in the regions Eq. (11) and obtain
$m$ $=$ $(15800\pm 90)~\mathrm{MeV},$
$\Lambda$ $=$ $(3.33\pm 0.35)~\mathrm{GeV}^{5}.$ (12)
The results in Eq. (12) are effectively equal to SR predictions at $M^{2}=17.5~\mathrm{GeV}^{2}$ and $s_{0}=280.5~\mathrm{GeV}^{2}$, where $\mathrm{PC}\approx 0.57$ ensuring the dominance of $\mathrm{PC}$ in $m$ and $\Lambda$. The uncertainties in Eq. (12) are equal to $\pm 0.6\%$ for the mass $m$, and to $\pm 11\%$ for the current coupling $\Lambda$. They appear mainly due to choices of $M^{2}$ and $s_{0}$: Ambiguities generated by errors in quark masses and gluon condensate are negligibly small. Note that these theoretical errors remain within limits accepted in SR analyses.
Dependence of $m$ on the parameters $M^{2}$ and $s_{0}$ is shown in Fig. 3. Having inspected curves presented there one can be convinced in stability of this prediction.
Figure 3: The mass $m$ as a function of $M^{2}$ (left panel), and $s_{0}$ (right panel).
Investigations aimed to find the parameters $\widetilde{m}$ and $\widetilde{\Lambda}$ of the hadronic molecule $\widetilde{\mathcal{M}}_{\mathrm{AV}}$ are carried out in accordance with a scheme described above. The QCD side of relevant SRs are given by the expression
[…]
$\left.\times S_{b}^{ba^{\prime}}(x)\gamma_{5}S_{b}^{a^{\prime}a}(-x)\right]-\mathrm{Tr}\left[\gamma_{5}S_{b}^{aa^{\prime}}(x)\gamma_{5}S_{b}^{a^{\prime}a}(-x)\right]$
$\left.\times\mathrm{Tr}\left[\gamma_{\mu}S_{b}^{bb^{\prime}}(x)\gamma_{\nu}S_{c}^{b^{\prime}b}(-x)\right]\right\},$ (13)
whereas their phenomenological component can be obtained from Eq. (6) by simple substitutions $m$, $\Lambda\rightarrow\widetilde{m}$, $\widetilde{\Lambda}$. Numerical computations confirm that SRs lead to almost identical predictions for $\widetilde{m}$, $\widetilde{\Lambda}$: Observed deviation, for instance, of m~\ $\widetilde${m} from $m$ amounts approximately to 12MeV1-2\ $\mathrm${MeV} which is beyond the accuracy of the used SR method. Therefore, below we consider decays of the hadronic molecule $\mathcal{M}_{\mathrm{AV}}$ and assume that $\widetilde{\mathcal{M}}_{\mathrm{AV}}$ has the same full width as $\mathcal{M}_{\mathrm{AV}}$.
III Decay of $\mathcal{M}_{\mathrm{AV}}$ to final states $\Upsilon B_{c}^{-}$ and $\eta_{b}B_{c}^{\ast-}$
Here, we study dominant decays $\mathcal{M}_{\mathrm{AV}}\rightarrow\Upsilon B_{c}^{-}$ and $\mathcal{M}_{\mathrm{AV}}\rightarrow\eta_{b}B_{c}^{\ast-}$ of the hadronic molecule $\mathcal{M}_{\mathrm{AV}}$, and calculate their partial widths. The result for the mass $m=(15800\pm 90)~\mathrm{MeV}$ of the molecule $\mathcal{M}_{\mathrm{AV}}$ shows that it exceeds the two-meson thresholds for production of the pairs $\Upsilon B_{c}^{-}$ and $\eta_{b}B_{c}^{\ast-}$. Indeed, the masses $m_{\Upsilon}=(9460.40\pm 0.10)~\mathrm{MeV}$ and $m_{B_{c}}=(6274.47\pm 0.27\pm 0.17)~\mathrm{MeV}$ of the mesons $\Upsilon$ and $B_{c}^{-}$, as well as $m_{\eta_{b}}=(9398.7\pm 2.0)~\mathrm{MeV}$ and $m_{B_{c}^{\ast}}=6338~\mathrm{MeV}$ of the mesons $\eta_{b}$ and $B_{c}^{\ast-}$ determine limits $15735~\mathrm{MeV}$ and $15737~\mathrm{MeV}$ which make possible dissociation to these particles. Note that $m_{\Upsilon}$, $m_{B_{c}}$, and $m_{\eta_{b}}$ are experimental data PDG:2024 , whereas for $m_{B_{c}^{\ast}}$ we use the model prediction Godfrey:2004ya .
It is worth emphasizing that $\mathcal{M}_{\mathrm{AV}}$ decays through these channels provided one employs for $m$ its central (or larger) value. In the lower limit of the mass $m$ decays to pairs $\Upsilon B_{c}^{-}$ and $\eta_{b}B_{c}^{\ast-}$ are kinematically forbidden processes. Then to estimate the full width of $\mathcal{M}_{\mathrm{AV}}$ one can consider only subleading channels. This problem will be addressed later in this work, but here we utilize as the mass of $\mathcal{M}_{\mathrm{AV}}$ the central value in our prediction.
III.1 Decay $\mathcal{M}_{\mathrm{AV}}\rightarrow\Upsilon B_{c}^{-}$
The width of the process $\mathcal{M}_{\mathrm{AV}}\rightarrow\Upsilon B_{c}^{-}$ contains, apart from other input parameters, also the strong coupling $g_{1}$ at the vertex $\mathcal{M}_{\mathrm{AV}}\Upsilon B_{c}^{-}$. The coupling $g_{1}$ can be extracted at the mass shell $q^{2}=m_{B_{c}}^{2}$ of the form factor $g_{1}(q^{2})$. The latter is evaluated using the SR obtained from analysis of the three-point correlator
$\Pi_{\mu\nu}^{1}(p,p^{\prime})$ $=$ $i^{2}\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}\langle 0|\mathcal{T}\{J_{\mu}^{\Upsilon}(y)$ (14)
$\times J^{B_{c}^{-}}(0)J_{\nu}^{\dagger}(x)\}|0\rangle,$
where $J_{\mu}^{\Upsilon}(x)$ and $J^{B_{c}^{-}}(x)$ are the interpolating currents of $\Upsilon$ and $B_{c}^{-}$, respectively. These currents are given by the expressions
$J_{\mu}^{\Upsilon}(x)=\overline{b}_{i}(x)\gamma_{\mu}b_{i}(x),\ J^{B_{c}^{-}}(x)=\overline{c}_{j}(x)i\gamma_{5}b_{j}(x).$ (15)
The four-momentum $p$ of the molecule $\mathcal{M}_{\mathrm{AV}}$ is connected to momenta $p^{\prime}$ and $q$ of produced particles, i.e., $p=p^{\prime}+q$.
The correlation function Eq. (14) expressed using parameters of the particles $\mathcal{M}_{\mathrm{AV}}$, $\Upsilon$ and $B_{c}^{-}$ is the physical component $\Pi_{\mu\nu}^{1\mathrm{Phys}}(p,p^{\prime})$ of the desired SR. To determine this correlator, we implement into Eq. (14) systems of intermediate states for these particles and perform four-dimensional integrations over $x$ and $y$. After separating the contribution of the ground-level states, we get
$\Pi_{\mu\nu}^{1\mathrm{Phys}}(p,p^{\prime})=\frac{\langle 0|J_{\mu}^{\Upsilon}|\Upsilon(p^{\prime},\varepsilon)\rangle}{p^{\prime 2}-m_{\Upsilon}^{2}}\frac{\langle 0|J^{B_{c}^{-}}|B_{c}^{-}(q)\rangle}{q^{2}-m_{B_{c}}^{2}}$
$\times\langle\Upsilon(p^{\prime},\varepsilon)B_{c}^{-}(q)|\mathcal{M}_{\mathrm{AV}}(p,\epsilon)\rangle\frac{\langle\mathcal{M}_{\mathrm{AV}}(p,\epsilon)|J_{\nu}^{\dagger}|0\rangle}{p^{2}-m^{2}}$
$+\cdots,$ (16)
with $\varepsilon_{\mu}$ being the polarization vector of $\Upsilon$. In Eq. (16) the dots encode contributions of excited and continuum states.
Having employed the matrix elements of the particles $\Upsilon$ and $B_{c}^{-}$
$\langle 0|J_{\mu}^{\Upsilon}|\Upsilon(p^{\prime},\varepsilon)\rangle$ $=$ $f_{\Upsilon}m_{\Upsilon}\varepsilon_{\mu},$
$\langle 0|J^{B_{c}^{-}}|B_{c}^{-}(q)\rangle$ $=$ $\frac{f_{B_{c}}m_{B_{c}}^{2}}{m_{b}+m_{c}},$ (17)
we rewrite $\Pi_{\mu\nu}^{1\mathrm{Phys}}(p,p^{\prime})$ in a form suitable for our purposes. In Eq. (17) $f_{\Upsilon}$ and $f_{B_{c}}$ are the decay constants of $\Upsilon$ and $B_{c}^{-}$, respectively. We also employ an expression for the vertex $\langle\Upsilon(p^{\prime},\varepsilon)B_{c}^{-}(q)|\mathcal{M}_{\mathrm{AV}}(p,\epsilon)\rangle$
$\langle\Upsilon(p^{\prime},\varepsilon)B_{c}^{-}(q)|\mathcal{M}_{\mathrm{AV}}(p,\epsilon)\rangle=g_{1}(q^{2})\left[(p\cdot q)(\epsilon\cdot\varepsilon^{\ast})\right$.
$\left.-(q\cdot\epsilon)(p\cdot\varepsilon^{\ast})\right].$ (18)
By applying these matrix elements, it is not difficult to find that
$\Pi_{\mu\nu}^{1\mathrm{Phys}}(p,p^{\prime})=g_{1}(q^{2})\frac{\Lambda f_{B_{c}}m_{B_{c}}^{2}f_{\Upsilon}m_{\Upsilon}}{(m_{b}+m_{c})\left(p^{2}-m^{2}\right)\left(p^{\prime 2}-m_{\Upsilon}^{2}\right)}$
$\times\frac{1}{(q^{2}-m_{B_{c}}^{2})}\left[\frac{m^{2}-m_{\Upsilon}^{2}+q^{2}}{2}g_{\mu\nu}-p_{\mu}p_{\nu}+p_{\mu}^{\prime}p_{\nu}\right$.
$\left.-\frac{m^{2}}{m_{\Upsilon}^{2}}p_{\mu}^{\prime}p_{\nu}^{\prime}+\frac{m^{2}+m_{\Upsilon}^{2}-q^{2}}{m_{\Upsilon}^{2}}p_{\mu}^{\prime}p_{\nu}\right]+\cdots.$ (19)
The function $\Pi_{\mu\nu}^{1\mathrm{Phys}}(p,p^{\prime})$ is a sum of two different Lorentz terms, one of which has be chosen for investigations. We continue with the invariant amplitude $\Pi_{1}^{\mathrm{Phys}}(p^{2},p^{\prime 2},q^{2})$ that corresponds to the structure $g_{\mu\nu}$.
The correlation function $\Pi_{\mu\nu}^{1}(p,p^{\prime})$ in terms of the quark propagators is equal to
[…]
$\left.\times\gamma_{\nu}S_{b}^{ai}(x-y)\right]\mathrm{Tr}\left[\gamma_{5}S_{b}^{jb}(-x)\gamma_{5}S_{c}^{bj}(x)\right]$
$\left.-\mathrm{Tr}\left[\gamma_{\mu}S_{b}^{ib}(y-x)\gamma_{5}S_{c}^{bj}(x)\gamma_{5}S_{b}^{ja}(-x)\gamma_{\nu}S_{b}^{ai}(x-y)\right]\right\}.$
We label by $\Pi_{1}^{\mathrm{OPE}}(p^{2},p^{\prime 2},q^{2})$ the invariant amplitude associated with the structure $\sim g_{\mu\nu}$ and utilize it in our studies.
By equating $\Pi_{1}^{\mathrm{Phys}}(p^{2},p^{\prime 2},q^{2})$ and $\Pi_{1}^{\mathrm{OPE}}(p^{2},p^{\prime 2},q^{2})$ and carrying out all required technical manipulations (Borel transformations over the variables $p^{2}$, $p^{\prime 2}$, subtraction of excited and continuum states’ contributions ), we determine SR for the form factor $g_{1}(q^{2})$
$g_{1}(q^{2})=\frac{2(m_{b}+m_{c})}{\Lambda f_{B_{c}}m_{B_{c}}^{2}f_{\Upsilon}m_{\Upsilon}}\frac{q^{2}-m_{B_{c}}^{2}}{m^{2}-m_{\Upsilon}^{2}+q^{2}}$
$\times e^{m^{2}/M_{1}^{2}}e^{m_{\Upsilon}^{2}/M_{2}^{2}}\Pi_{1}(\mathbf{M}^{2},\mathbf{s}_{0},q^{2}),$ (21)
where
$\Pi_{1}(\mathbf{M}^{2},\mathbf{s}_{0},q^{2})=\int_{(3m_{b}+m_{c})^{2}}^{s_{0}}ds\int_{4m_{b}^{2}}^{s_{0}^{\prime}}ds^{\prime}\rho_{1}(s,s^{\prime},q^{2})$
$\times e^{-s/M_{1}^{2}}e^{-s^{\prime}/M_{2}^{2}}.$ (22)
In Eqs. (21) and (22) $\Pi_{1}(\mathbf{M}^{2},\mathbf{s}_{0},q^{2})$ is the amplitude $\Pi_{1}^{\mathrm{OPE}}(p^{2},p^{\prime 2},q^{2})$ after the double Borel and continuum subtraction procedures. The spectral density $\rho(s,s^{\prime},q^{2})$ in the expression above is equal to the imaginary part of $\Pi_{1}^{\mathrm{OPE}}(s,s^{\prime},q^{2})$.
The function $\Pi_{1}(\mathbf{M}^{2},\mathbf{s}_{0},q^{2})$ contains parameters $\mathbf{M}^{2}=(M_{1}^{2},M_{2}^{2})$ and $\mathbf{s}_{0}=(s_{0},s_{0}^{\prime})$ where $(M_{1}^{2},s_{0})$ and $(M_{2}^{2},s_{0}^{\prime})$ correspond to channels of the particles $\mathcal{M}_{\mathrm{AV}}$ and $\Upsilon$. They should be constrained in accordance with usual rules of SR computations which have been explained in the previous section. Numerical analysis proves that Eq. (11) for the parameters $(M_{1}^{2},s_{0})$ and
$M_{2}^{2}\in[10,12]~\mathrm{GeV}^{2},\ s_{0}^{\prime}\in[98,100]~\mathrm{GeV}^{2}.$ (23)
for $(M_{2}^{2},s_{0}^{\prime})$ satisfy these constraints.
For numerical computations one needs also the spectroscopic parameters of the mesons $\Upsilon$ and $B_{c}^{-}$. There are experimental information on masses of these particles PDG:2024 . Their decay constants $f_{\Upsilon}=(708\pm 8)~\mathrm{MeV}$ and $f_{B_{c}}=(371\pm 37)~\mathrm{MeV}$ are borrowed from Refs. Lakhina:2006vg ; Wang:2024fwc , respectively. The SR approach generates reliable predictions in the Euclidean region $q^{2}<0$, whereas $g_{1}(q^{2})$ becomes equal to $g_{1}$ at the mass shell $q^{2}=m_{B_{c}}^{2}$. Therefore, it is convenient to introduce the new function $g_{1}(Q^{2})$, where $Q^{2}=-q^{2}$, and apply it in future studies. The QCD predictions for $g_{1}(Q^{2})$ are depicted in Fig. 4, where $Q^{2}$ changes inside limits $Q^{2}=2-30~\mathrm{GeV}^{2}$.
To extract $g_{1}$ at $q^{2}=-Q^{2}=m_{B_{c}}^{2}$, we introduce the fit function $\mathcal{F}_{1}(Q^{2})$: For $Q^{2}>0$ it amounts to the QCD data, but can also be extended to the region of $Q^{2}<0$. We choose this function in the following form
$\mathcal{F}_{i}(Q^{2})=\mathcal{F}_{i}^{0}\mathrm{\exp}\left[l_{i}^{1}\frac{Q^{2}}{m^{2}}+l_{i}^{2}\left(\frac{Q^{2}}{m^{2}}\right)^{2}\right],$ (24)
where the parameters $\mathcal{F}_{i}^{0}$, $l_{i}^{1}$, and $l_{i}^{2}$ have to be extracted from comparison of $\mathcal{F}_{1}(Q^{2})$ and the QCD data. It is easy to find them:
$\mathcal{F}_{1}^{0}=3.21~\mathrm{GeV}^{-1},l_{1}^{1}=15.40,\text{and }l_{1}^{2}=-12.40.$ (25)
In Fig. 4 we also show $\mathcal{F}_{1}(Q^{2})$: A reasonable agreement of $\mathcal{F}_{1}(Q^{2})$ and QCD data is clear. As a result, for $g_{1}$ we find
$g_{1}\equiv\mathcal{F}_{1}(-m_{B_{c}}^{2})=(2.0\pm 0.4)\times 10^{-1}\ \mathrm{GeV}^{-1}.$ (26)
The width of the process $\mathcal{M}_{\mathrm{AV}}\rightarrow\Upsilon B_{c}^{-}$ can be calculated by means of the expression
$\Gamma\left[\mathcal{M}_{\mathrm{AV}}\rightarrow\Upsilon B_{c}^{-}\right]=g_{1}^{2}\frac{\lambda_{1}}{24\pi m^{2}}|M_{1}|^{2},$ (27)
where
$|M_{1}|^{2}$ $=$ $\frac{1}{4m_{\Upsilon}^{2}}\left[m^{6}-2m^{4}m_{B_{c}}^{2}+2m_{\Upsilon}^{2}(m_{\Upsilon}^{2}-m_{B_{c}}^{2})^{2}\right$. (28)
$\left.+m^{2}(m_{B_{c}}^{4}+6m_{B_{c}}^{2}m_{\Upsilon}^{2}-3m_{\Upsilon}^{4})\right].$
We have used also the parameter $\lambda_{1}=\lambda(m,m_{\Upsilon},m_{B_{c}})$ with $\lambda(a,b,c)$ being defined as
$\lambda(a,b,c)=\frac{\sqrt{a^{4}+b^{4}+c^{4}-2(a^{2}b^{2}+a^{2}c^{2}+b^{2}c^{2})}}{2a}.$ (29)
Finally, we get,
$\Gamma\left[\mathcal{M}_{\mathrm{AV}}\rightarrow\Upsilon B_{c}^{-}\right]=(46.9\pm 13.3)~\mathrm{MeV}.$ (30)
Figure 4: QCD data and extrapolating functions $\mathcal{F}_{1}(Q^{2})$ (solid curve) and $\mathcal{F}_{2}(Q^{2})$ (dashed line). The diamond and circle fix the points $Q^{2}=-m_{B_{c}}^{2}$ and $Q^{2}=-m_{B_{c}^{\ast}}^{2}$, respectively.
III.2 Process $\mathcal{M}_{\mathrm{AV}}\rightarrow\eta_{b}B_{c}^{\ast-}$
The process $\mathcal{M}_{\mathrm{AV}}\rightarrow\eta_{b}B_{c}^{\ast-}$ is studied in accordance with the scheme outlined above. Here, we are going to estimate the coupling $g_{2}$ at $\mathcal{M}_{\mathrm{AV}}\eta_{b}B_{c}^{\ast-}$. For these purposes, we start from analysis of the correlation function $\Pi_{\mu\nu}^{2}(p,p^{\prime})$ that should allow us to evaluate the form factor $g_{2}(q^{2})$: The latter at the mass shell $q^{2}=m_{B_{c}^{\ast}}^{2}$ gives $g_{2}$.
This correlation function is introduced by means of the formula
$\Pi_{\mu\nu}^{2}(p,p^{\prime})$ $=$ $i^{2}\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}\langle 0|\mathcal{T}\{J^{\eta_{b}}(y)$ (31)
$\times J_{\mu}^{B_{c}^{\ast}}(0)J_{\nu}^{\dagger}(x)\}|0\rangle,$
with $J^{\eta_{b}}(x)$ and $J_{\mu}^{B_{c}^{\ast}}(x)$ being the currents that interpolate particles $\eta_{b}$ and $B_{c}^{\ast-}$. They have the following forms
$J^{\eta_{b}}(x)=\overline{b}_{i}(x)i\gamma_{5}b_{i}(x),\ J_{\mu}^{B_{c}^{\ast}}(x)=\overline{c}_{j}(x)\gamma_{\mu}b_{j}(x).$ (32)
The correlator $\Pi_{\mu\nu}^{2}(p,p^{\prime})$ in terms of physical parameters of the particles $\mathcal{M}_{\mathrm{AV}}$, $\eta_{b}$, and $B_{c}^{\ast-}$ is given by the expression
$\Pi_{\mu\nu}^{2\mathrm{Phys}}(p,p^{\prime})=\frac{g_{2}(q^{2})\Lambda f_{\eta_{b}}m_{\eta_{b}}^{2}f_{B_{c}^{\ast}}m_{B_{c}^{\ast}}}{2m_{b}\left(p^{2}-m^{2}\right)\left(p^{\prime 2}-m_{\eta_{b}}^{2}\right)(q^{2}-m_{B_{c}^{\ast}}^{2})}$
$\times\left[\frac{m^{2}+m_{\eta_{b}}^{2}-q^{2}}{2}g_{\mu\nu}-\frac{m^{2}}{m_{B_{c}^{\ast}}^{2}}p_{\mu}^{\prime}p_{\nu}^{\prime}+\frac{m^{2}-m_{B_{c}^{\ast}}^{2}}{m_{B_{c}^{\ast}}^{2}}p_{\mu}^{\prime}p_{\nu}\right$.
$\left.-\frac{m^{2}+m_{\eta_{b}}^{2}-q^{2}}{2m_{B_{c}^{\ast}}^{2}}(p_{\mu}p_{\nu}+p_{\mu}p_{\nu}^{\prime})\right]+\cdots.$ (33)
To derive $\Pi_{\mu\nu}^{2\mathrm{Phys}}(p,p^{\prime})$ we have employed the matrix elements
$\langle 0|J_{\mu}^{B_{c}^{\ast}}|B_{c}^{\ast-}(q,\varepsilon)\rangle$ $=$ $f_{B_{c}^{\ast}}m_{B_{c}^{\ast}}\varepsilon_{\mu},$
$\langle 0|J^{\eta_{b}}|\eta_{b}(p^{\prime})\rangle$ $=$ $\frac{f_{\eta_{b}}m_{\eta_{b}}^{2}}{2m_{b}},$ (34)
and
$\langle\eta_{b}(p^{\prime})B_{c}^{\ast-}(q,\varepsilon)|\mathcal{M}_{\mathrm{AV}}(p,\epsilon)\rangle=g_{2}(q^{2})\left[(p\cdot p^{\prime})(\epsilon\cdot\varepsilon^{\ast})\right$.
$\left.-(p^{\prime}\cdot\epsilon)(p\cdot\varepsilon^{\ast}\right].$ (35)
The QCD side of the SR is determined by the formula
$\Pi_{\mu\nu}^{2\mathrm{OPE}}(p,p^{\prime})=-\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}\mathrm{Tr}\left[\gamma_{5}S_{b}^{ib}(y-x)\right$.
$\left.\times\gamma_{5}S_{c}^{bj}(x)\gamma_{\mu}S_{b}^{ja}(-x)\gamma_{\nu}S_{b}^{ai}(x-y)\right].$ (36)
To find SR for the form factor $g_{2}(q^{2})$ we use the invariant amplitudes $\Pi_{2}^{\mathrm{Phys}}(p^{2},p^{\prime 2},q^{2})$ and $\Pi_{2}^{\mathrm{OPE}}(p^{2},p^{\prime 2},q^{2})$ related to structures $\sim g_{\mu\nu}$ both in the phenomenological and $\mathrm{OPE}$ versions of the correlation function $\Pi_{\mu\nu}^{2}(p,p^{\prime})$. Then, the required SR reads
$g_{2}(q^{2})=\frac{4m_{b}}{\Lambda f_{\eta_{b}}m_{\eta_{b}}^{2}f_{B_{c}^{\ast}}m_{B_{c}^{\ast}}}\frac{q^{2}-m_{B_{c}^{\ast}}^{2}}{m^{2}+m_{\eta_{b}}^{2}-q^{2}}$
$\times e^{m^{2}/M_{1}^{2}}e^{m_{\eta_{b}}^{2}/M_{2}^{2}}\Pi_{2}(\mathbf{M}^{2},\mathbf{s}_{0},q^{2}).$ (37)
Numerical computations are carried out by employing the decay constants $f_{\eta_{b}}=724~\mathrm{MeV}$ and $f_{B_{c}^{\ast}}=471~\mathrm{MeV}$ Eichten:2019gig , respectively. The regions
$M_{2}^{2}\in[10,12]~\mathrm{GeV}^{2},\ s_{0}^{\prime}\in[95,99]~\mathrm{GeV}^{2},$ (38)
in the channel of $\eta_{b}$ meson meet necessary restrictions of SR analysis. The relevant extrapolating function $\mathcal{F}_{2}(Q^{2})$ has the parameters
$\mathcal{F}_{2}^{0}=0.433~\mathrm{GeV}^{-1},l_{2}^{1}=4.47,\text{and }l_{2}^{2}=-9.31.$ (39)
The SR data and the function $\mathcal{F}_{2}(Q^{2})$ are plotted in Fig. 4 as well. We obtain for $g_{2}$
$g_{2}\equiv\mathcal{F}_{2}(-m_{B_{c}}^{2})=(1.7\pm 0.3)\times 10^{-1}\ \mathrm{GeV}^{-1}.$ (40)
To evaluate the width of this decay we employ the following expression
$\Gamma\left[\mathcal{M}_{\mathrm{AV}}\rightarrow\eta_{b}B_{c}^{\ast-}\right]=g_{2}^{2}\frac{\lambda_{2}}{24\pi m^{2}}|M_{2}|^{2},$ (41)
where $|M_{2}|$ is obtainable from $|M_{1}|$ upon $m_{B_{c}}$, $m_{\Upsilon}\rightarrow m_{\eta_{b}}$, $m_{B_{c}^{\ast}}$ substitutions. Now $\lambda_{2}$ amounts to $\lambda(m,m_{\eta_{b}},m_{B_{c}^{\ast}})$. The width of this process is
$\Gamma\left[\mathcal{M}_{\mathrm{AV}}\rightarrow\eta_{b}B_{c}^{\ast-}\right]=(31.7\pm 9.5)~\mathrm{MeV}.$ (42)
IV Decays generated by annihilation of $\overline{b}b$ quarks
Annihilation of $\overline{b}b$ quarks triggers various decay modes of the hadronic molecular state $\mathcal{M}_{\mathrm{AV}}$. Decays $\mathcal{M}_{\mathrm{AV}}\rightarrow B^{\ast-}\overline{D}^{0}$, $\overline{B}^{\ast 0}D^{-}$, $B^{-}\overline{D}^{\ast 0}$, $\overline{B}^{0}D^{\ast-}$, $\overline{B}_{s}^{\ast 0}D_{s}^{-}$, and $\overline{B}_{s}^{0}D_{s}^{\ast-}$ are such channels. Note that parameters some of these processes are very close to each other. This is connected with the following fact: In this work we use the approximation $m_{\mathrm{u}}=m_{\mathrm{d}}=0$ and $m_{\mathrm{s}}=(93.5\pm 0.8)~\mathrm{MeV}$. Then it is not difficult to see that the QCD sides of the SRs for the channels $\mathcal{M}_{\mathrm{AV}}\rightarrow B^{\ast-}\overline{D}^{0}$ and $\overline{B}^{\ast 0}D^{-}$, as well as for $\mathcal{M}_{\mathrm{AV}}\rightarrow B^{-}\overline{D}^{\ast 0}$and $\overline{B}^{0}D^{\ast-}$ are given by the same expressions. Differences in the physical sides are connected with the masses of mesons, for example, $\overline{D}^{0}$ and $D^{-}$ which are small. Therefore, we consider these processes as identical channels including them into two groups and calculate the partial width one of them.
IV.1 $\mathcal{M}_{\mathrm{AV}}\rightarrow B^{\ast-}\overline{D}^{0}$, $\overline{B}^{\ast 0}D^{-}$
In this subsection, we investigate the channel $\mathcal{M}_{\mathrm{AV}}\rightarrow B^{\ast-}\overline{D}^{0}$ of the molecule $\mathcal{M}_{\mathrm{AV}}$ and compute its partial decay width. We are going to estimate the coupling $G$ at the vertex $\mathcal{M}_{\mathrm{AV}}B^{\ast-}\overline{D}^{0}$. The correlation function suitable for our purposes is
$\overline{\Pi}_{\mu\nu}(p,p^{\prime})$ $=$ $i^{2}\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}\langle 0|\mathcal{T}\{J_{\mu}^{B^{\ast}}(y)$ (43)
$\times J^{\overline{D}^{0}}(0)J_{\nu}^{\dagger}(x)\}|0\rangle,$
where $J_{\mu}^{B^{\ast}}$ and $J^{\overline{D}^{0}}(x)$ are currents for $B^{\ast-}$ and $\overline{D}^{0}$ mesons which are given by the expressions
$J_{\mu}^{B^{\ast}}(x)=\overline{u}_{i}(x)\gamma_{\mu}b_{i}(x),\ J^{\overline{D}^{0}}(x)=\overline{c}_{j}(x)i\gamma_{5}u_{j}(x).$ (44)
The phenomenological expression for $\overline{\Pi}_{\mu\nu}(p,p^{\prime})$ is obtained using the matrix elements
$\langle 0|J_{\mu}^{B^{\ast}}|B^{\ast-}(p^{\prime},\varepsilon)\rangle$ $=$ $f_{B^{\ast}}m_{B^{\ast}}\varepsilon_{\mu},$
$\langle 0|J^{\overline{D}^{0}}|\overline{D}^{0}(q)\rangle$ $=$ $\frac{f_{D}m_{\overline{D}^{0}}^{2}}{m_{c}}.$ (45)
In Eq. (45) $m_{B^{\ast}}=(5324.75\pm 0.20)~\mathrm{MeV},\ m_{\overline{D}^{0}}=(1864.84\pm 0.05)~\mathrm{MeV}$ and $f_{B^{\ast}}=(210\pm 6)~\mathrm{MeV},$ $f_{D}=(211.9\pm 1.1)~\mathrm{MeV}$ are the masses and decay constants of the these mesons, whereas $\varepsilon_{\mu}$ is the polarization vector of the vector meson $B^{\ast-}$. The vertex $\langle B^{\ast-}(p^{\prime},\varepsilon)\overline{D}^{0}(q)|B^{\ast-}(p,\epsilon)\rangle$ and correlator $\overline{\Pi}_{\mu\nu}^{\mathrm{Phys}}(p,p^{\prime})$ are analogous to ones presented in section III.
Calculations of $\overline{\Pi}_{\mu\nu}(p,p^{\prime})$ using the quark propagators yield
$\overline{\Pi}_{\mu\nu}^{\mathrm{OPE}}(p,p^{\prime})=\frac{\langle\overline{b}b\rangle}{3}\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}\mathrm{Tr}\left[\gamma_{\mu}S_{b}^{ia}(y-x)\right$.
$\left.\times\gamma_{\nu}\gamma_{5}S_{c}^{aj}(x)\gamma_{5}S_{u}^{ji}(-y)\right].$ (46)
We use also the relation
$\langle\overline{b}b\rangle=-\frac{1}{12m_{b}}\langle\frac{\alpha_{s}G^{2}}{\pi}\rangle$ (47)
obtained in Ref. Shifman:1978bx in the framework of SR method.
The form factor $G(Q^{2})$ that at $Q^{2}=-m_{\overline{D}^{0}}^{2}$ amounts to $G$ is computed in the region $Q^{2}=2-30\ \mathrm{GeV}^{2}$. For $(M_{1}^{2},s_{0})$ we have used their values from Eq. (11), and the parameters $(M_{2}^{2},s_{0}^{\prime})$ have been varied within the borders
$M_{2}^{2}\in[5.5,6.5]~\mathrm{GeV}^{2},\ s_{0}^{\prime}\in[34,35]~\mathrm{GeV}^{2}.$ (48)
Results for $G(Q^{2})$ are plotted in Fig. 5. The extrapolating function $\overline{\mathcal{F}}(Q^{2})$ is determined by the constants $\overline{\mathcal{F}}^{0}=0.0166~\mathrm{GeV}^{-1}$, $\overline{l}^{1}=13.28$, and $\overline{l}^{2}=-16.70.$ The strong coupling $G$ is evaluated at the mass shell $q^{2}=m_{\overline{D}^{0}}^{2}$ and equal to
$G\equiv\overline{\mathcal{F}}(-m_{\overline{D}^{0}}^{2})=(1.37\pm 0.31)\times 10^{-2}~\mathrm{GeV}^{-1}.$ (49)
For the width of the process $\mathcal{M}_{\mathrm{AV}}\rightarrow B^{\ast-}\overline{D}^{0}$ our analysis predicts
$\Gamma\left[\mathcal{M}_{\mathrm{AV}}\rightarrow B^{\ast-}\overline{D}^{0}\right]=(8.9\pm 2.9)~\mathrm{MeV}.$ (50)
Let us note that ambiguities above are total errors generated by uncertainties in $G$ and in the masses of the particles $\mathcal{M}_{\mathrm{AV}}$, $B^{\ast-}$, and $\overline{D}^{0}$.
The width of the second process $\mathcal{M}_{\mathrm{AV}}\rightarrow\overline{B}^{\ast 0}D^{-}$, as it has been discussed above, is approximately equal to that of the first decay. In other words
$\Gamma\left[\mathcal{M}_{\mathrm{AV}}\rightarrow\overline{B}^{\ast 0}D^{-}\right]\approx\Gamma\left[\mathcal{M}_{\mathrm{AV}}\rightarrow B^{\ast-}\overline{D}^{0}\right].$ (51)
Figure 5: SR data for the form factors $G(Q^{2})$ and $\widehat{G}_{1}(Q^{2})$ and fit functions $\overline{\mathcal{F}}(Q^{2})$ (solid line), $\widehat{\mathcal{F}}_{1}(Q^{2})$ (dash-dotted line). The circle and star are placed at positions $Q^{2}=-m_{\overline{D}^{0}}^{2}$ and $Q^{2}=-m_{D_{s}^{-}}^{2}$, respectively.
IV.2 $\mathcal{M}_{\mathrm{AV}}\rightarrow B^{-}\overline{D}^{\ast 0}$, $\overline{B}^{0}D^{\ast-}$
These channels are studied within the same approach. Here, we give formulas for the decay $\mathcal{M}_{\mathrm{AV}}\rightarrow B^{-}\overline{D}^{\ast 0}$ and our estimate its partial width: The partial width of the second mode $\mathcal{M}_{\mathrm{AV}}\rightarrow\overline{B}^{0}D^{\ast-}$ is approximately equal to parameters of the first decay.
We start from studying the correlation function
$\widetilde{\Pi}_{\mu\nu}(p,p^{\prime})$ $=$ $i^{2}\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}\langle 0|\mathcal{T}\{J^{B}(y)$ (52)
$\times J_{\mu}^{\overline{D}^{\ast 0}}(0)J_{\nu}^{\dagger}(x)\}|0\rangle.$
Here, $J^{B}(x)$ and $J_{\mu}^{\overline{D}^{\ast 0}}(x)$ are relevant interpolating currents
$J^{B}(x)=\overline{u}_{i}(x)i\gamma_{5}b_{i}(x),\ J_{\mu}^{\overline{D}^{\ast 0}}(x)=\overline{c}_{j}(x)\gamma_{\mu}u_{j}(x).$ (53)
The correlator in Eq. (52) is required to find SR for the form factor $\widetilde{G}(q^{2})$ that at the mass shell of the meson $\overline{D}^{\ast 0}$ permits us to evaluate the strong coupling $\widetilde{G}$ at the vertex $\mathcal{M}_{\mathrm{AV}}\overline{B}^{0}D^{\ast-}$.
We calculate the correlator $\widetilde{\Pi}_{\mu\nu}^{\mathrm{Phys}}(p,p^{\prime})$ by employing the following matrix elements
$\langle 0|J^{B}|B^{-}(p^{\prime})\rangle$ $=$ $\frac{f_{B}m_{B}^{2}}{m_{b}},$
$\langle 0|J_{\mu}^{\overline{D}^{\ast 0}}|\overline{D}^{\ast 0}(q,\varepsilon)\rangle$ $=$ $f_{\overline{D}^{\ast 0}}m_{\overline{D}^{\ast 0}}\varepsilon_{\mu},$ (54)
where $m_{B}=(5279.41\pm 0.07)~\mathrm{MeV}$, $f_{B}=206~\mathrm{MeV}$ and $m_{\overline{D}^{\ast 0}}=(2006.85\pm 0.05)~\mathrm{MeV}$, $f_{\overline{D}^{\ast 0}}=(252.2\pm 22.66)~\mathrm{MeV}$ are spectroscopic parameters of the mesons under discussion. The polarization vector of the $\overline{D}^{\ast 0}$ meson is labeled by $\varepsilon_{\mu}$.
The QCD component of SR is found in the form
$\widetilde{\Pi}_{\mu\nu}^{\mathrm{OPE}}(p,p^{\prime})=\frac{\langle\overline{b}b\rangle}{3}\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}\mathrm{Tr}\left[\gamma_{5}S_{b}^{ia}(y-x)\right$.
$\left.\times\gamma_{\nu}\gamma_{5}S_{c}^{aj}(x)\gamma_{\mu}S_{u}^{ji}(-y)\right].$ (55)
Numerical computations of SR for $\widetilde{G}(Q^{2})$ fulfilled at $Q^{2}=2-30\ \mathrm{GeV}^{2}$ using the regions Eq. (11) for $(M_{1}^{2},s_{0})$ and
$M_{2}^{2}\in[5.5,6.5]~\mathrm{GeV}^{2},\ s_{0}^{\prime}\in[33.5,34.5]~\mathrm{GeV}^{2},$ (56)
for the parameters $(M_{2}^{2},s_{0}^{\prime})$. The fit function $\widetilde{\mathcal{F}}(Q^{2})$ that employed to fix the coupling $\widetilde{G}$ has the following parameters $\widetilde{\mathcal{F}}^{0}=0.0049~\mathrm{GeV}^{-1},\widetilde{l}^{1}=12.96,\widetilde{l}^{2}=-16.96.$ This function leads to the predictions
$\widetilde{G}\equiv\widetilde{\mathcal{F}}(-m_{\overline{D}^{\ast 0}}^{2})=(3.94\pm 0.78)\times 10^{-3}\ \mathrm{GeV}^{-1},$ (57)
and
$\Gamma\left[\mathcal{M}_{\mathrm{AV}}\rightarrow B^{-}\overline{D}^{\ast 0}\right]=(4.4\pm 1.3)~\mathrm{MeV}.$ (58)
The partial width of the process $\mathcal{M}_{\mathrm{AV}}\rightarrow\overline{B}^{0}D^{\ast-}$ also amounts to Eq. (58).
IV.3 $\mathcal{M}_{\mathrm{AV}}\rightarrow\overline{B}_{s}^{\ast 0}D_{s}^{-}$, $\overline{B}_{s}^{0}D_{s}^{\ast-}$
Remaining two modes of $\mathcal{M}_{\mathrm{AV}}$ for investigations are its decays to the pairs of mesons $\overline{B}_{s}^{\ast 0}D_{s}^{-}$ and $\overline{B}_{s}^{0}D_{s}^{\ast-}$. We consider in detail the channel $\mathcal{M}_{\mathrm{AV}}\rightarrow\overline{B}_{s}^{\ast 0}D_{s}^{-}$ and provide results for the second process.
For the decay $\mathcal{M}_{\mathrm{AV}}\rightarrow\overline{B}_{s}^{\ast 0}D_{s}^{-}$ one should analyze the correlation function
$\widehat{\Pi}_{\mu\nu}^{1}(p,p^{\prime})$ $=$ $i^{2}\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}\langle 0|\mathcal{T}\{J_{\mu}^{B_{s}^{\ast}}(y)$ (59)
$\times J^{D_{s}}(0)J_{\nu}^{\dagger}(x)\}|0\rangle,$
[…]
$J_{\mu}^{B_{s}^{\ast}}(x)=\overline{s}_{i}(x)\gamma_{\mu}b_{i}(x),~J^{D_{s}}(x)=\overline{c}_{j}(x)i\gamma_{5}s_{j}(x).$ (60)
To find the phenomenological component of SR we apply the matrix elements
$\langle 0|J_{\mu}^{B^{\ast}}|\overline{B}_{s}^{\ast 0}(p^{\prime},\varepsilon)\rangle=f_{B_{s}^{\ast}}m_{B_{s}^{\ast}}\varepsilon_{\mu},$
$\langle 0|J^{D_{s}}|D_{s}^{-}(q)\rangle=\frac{f_{D_{s}}m_{D_{s}}^{2}}{m_{c}+m_{s}}.$ (61)
The vertex $\langle\overline{B}^{\ast 0}(p^{\prime},\varepsilon)D_{s}^{-}(q)|\mathcal{M}_{\mathrm{AV}}(p,\epsilon)\rangle$ has the standard form. In Eq. (61) $m_{B_{s}^{\ast}}=(5415.4\pm 1.4)~\mathrm{MeV}$, $f_{B_{s}^{\ast}}=221~\mathrm{MeV}$ and $m_{D_{s}}=(1968.35\pm 0.07)~\mathrm{MeV}$, $f_{D_{s}}=(249.9\pm 0.5)~\mathrm{MeV}$ are the parameters of the final-state mesons.
The correlator $\widehat{\Pi}_{\mu\nu}^{1}(p,p^{\prime})$ in terms of the heavy quark propagators is
$\widehat{\Pi}_{\mu\nu}^{1\mathrm{OPE}}(p,p^{\prime})$ $=$ $\frac{\langle\overline{b}b\rangle}{3}\int d^{4}xd^{4}ye^{ip^{\prime}y}e^{-ipx}\mathrm{Tr}\left[\gamma_{\mu}S_{b}^{ia}(y-x)\right$. (62)
$\left.\times\gamma_{\nu}\gamma_{5}S_{c}^{aj}(x)\gamma_{5}S_{s}^{ji}(-y)\right].$
The SR for the form factor $\widehat{G}_{1}(Q^{2})$ is found by utilizing amplitudes connected with the terms $\sim g_{\mu\nu}$ both in $\Pi_{\mu\nu}^{\mathrm{Phys}}(p,p^{\prime})$ and $\Pi_{\mu\nu}^{\mathrm{OPE}}(p,p^{\prime})$.
Numerical calculations are performed by employing
$M_{2}^{2}\in[6,7]~\mathrm{GeV}^{2},\ s_{0}^{\prime}\in[35,36]~\mathrm{GeV}^{2},$ (63)
for $\overline{B}_{s}^{\ast 0}$ channel. The constants in the fit function $\widehat{\mathcal{F}}_{1}(Q^{2})$ are $\widehat{\mathcal{F}}_{1}^{0}=0.0119~\mathrm{GeV}^{-1}$, $\widehat{l}_{1}^{1}=13.11$, and $\widehat{l}_{1}^{2}=-16.36$. The corresponding strong coupling $\widehat{G}_{1}$ is equal to
$\widehat{G}_{1}=(9.75\pm 1.87)\times 10^{-3}\ \mathrm{GeV}^{-1}.$ (64)
Then, the partial width of the decay $\mathcal{M}_{\mathrm{AV}}\rightarrow\overline{B}_{s}^{\ast 0}D_{s}^{-}$ amounts to
$\Gamma\left[\mathcal{M}_{\mathrm{AV}}\rightarrow\overline{B}_{s}^{\ast 0}D_{s}^{-}\right]=(4.3\pm 1.2)~\mathrm{MeV}.$ (65)
The next channel with the final-state strange mesons $\mathcal{M}_{\mathrm{AV}}\rightarrow\overline{B}_{s}^{0}D_{s}^{\ast-}$ is treated in the context of similar analyses. Our prediction for the coupling $\widehat{G}_{2}$ reads
$\widehat{G}_{2}=(3.21\pm 0.67)\times 10^{-3}\ \mathrm{GeV}^{-1}.$ (66)
Then, the partial width of the decay $\mathcal{M}_{\mathrm{AV}}\rightarrow\overline{B}_{s}^{0}D_{s}^{\ast-}$ equals to
$\Gamma\left[\mathcal{M}_{\mathrm{AV}}\rightarrow\overline{B}_{s}^{0}D_{s}^{\ast-}\right]=(2.6\pm 0.8)~\mathrm{MeV}.$ (67)
Note that this result has been obtained utilizing the parameters
$M_{2}^{2}\in[5.5,6.5]~\mathrm{GeV}^{2},\ s_{0}^{\prime}\in[34,35]~\mathrm{GeV}^{2},$ (68)
in $\overline{B}_{s}^{0}$ meson’s channel. The coupling $\widehat{G}_{2}$ has been estimated by means of the extrapolating function $\widehat{\mathcal{F}}_{2}(Q^{2})$ with $\widehat{\mathcal{F}}_{2}^{0}=0.004~\mathrm{GeV}^{-1}$, $\widehat{l}_{2}^{1}=12.78$, and $\widehat{l}_{2}^{2}=-16.64$.
By taking into account all results obtained in this and previous sections, it is easy to evaluate the full decay width of the hadronic molecule $\mathcal{M}_{\mathrm{AV}}$:
$\Gamma\left[\mathcal{M}_{\mathrm{AV}}\right]=(112\pm 17)~\mathrm{MeV}.$ (69)
This result shows that $\mathcal{M}_{\mathrm{AV}}$ is a rather broad compound, where the subleading decays are essential and form up to $30\%$ of its full width.
V Analysis and Final Notes
In the present article, we have continued our studies of the exotic hadronic molecular structures with asymmetric quark contents. In Ref. Agaev:2025did , we explored the scalar states $\eta_{b}B_{c}^{-}$ and $\eta_{c}B_{c}^{-}$ with contents $bb\overline{b}\overline{c}$ and $cc\overline{c}\overline{b}$, respectively. In this work, we have considered the axial-vector structures $\mathcal{M}_{\mathrm{AV}}=\Upsilon B_{c}^{-}$ and $\widetilde{\mathcal{M}}_{\mathrm{AV}}=\eta_{b}B_{c}^{\ast-}$ by calculating their masses and decay widths. It turned out that these particles have, at least within accuracy of the SR method, identical masses. Therefore, we have assumed that their decay widths are approximately equal to each other and studied in detail decay channels of $\mathcal{M}_{\mathrm{AV}}$.
Our prediction $m=(15800\pm 90)~\mathrm{MeV}$ for the mass of the molecule $\mathcal{M}_{\mathrm{AV}}$ means that it is strong-interaction unstable particle and decays to $\Upsilon B_{c}^{-}$ and $\eta_{b}B_{c}^{\ast-}$ pairs of ordinary heavy mesons. Apart from these two dominant modes, there are channels of $\mathcal{M}_{\mathrm{AV}}$ which become possible due to annihilation of $b\overline{b}$ quarks in $\mathcal{M}_{\mathrm{AV}}$ and generation of light quarks’ pairs. We have investigated decays to $B^{\ast-}\overline{D}^{0}$, $\overline{B}^{\ast 0}D^{-}$, $B^{-}\overline{D}^{\ast 0}$$\overline{B}^{0}D^{\ast-}$, $\overline{B}_{s}^{\ast 0}D_{s}^{-}$, and $\overline{B}_{s}^{0}D_{s}^{\ast-}$ mesons from the class of subleading processes.
The fully heavy hadronic molecules were object of interesting analyses in Ref. Liu:2024pio , in which the authors applied the extended local gauge formalism to investigate such systems with different spin-parities. The masses of the hadronic molecules were estimated there using the cutoff momentum $\Lambda$ of the model. The authors made interesting conclusions about properties of the fully heavy axial-vector $\Upsilon B_{c}^{-}$, $\eta_{b}B_{c}^{\ast-}$ and tensor $\Upsilon B_{c}^{\ast-}$ molecules. In accordance with their predictions, these structures reside below the corresponding two-meson thresholds and form bound states. The mass of the axial-vector molecules $\Upsilon B_{c}^{-}$ and $\eta_{b}B_{c}^{\ast-}$ was found there equal to $15725.3~\mathrm{MeV}$ at $\Lambda=1000~\mathrm{MeV}$.
This result is smaller than our prediction provided one compares it with the central value $m=15800~\mathrm{MeV}$, but at the lower limit, $m=15710~\mathrm{MeV}$, they are comparable with each other. The hadronic molecule with the mass $15710~\mathrm{MeV}$, of course, can not decay to $\Upsilon B_{c}^{-}$ and $\eta_{b}B_{c}^{\ast-}$ mesons. Nevertheless the subleading channels for its transformation to conventional particles remain open is this case as well. To elucidate this question we have computed the width of $\mathcal{M}_{\mathrm{AV}}$ with the mass $15710~\mathrm{MeV}$. Analysis has been carried out in accordance with the scheme presented in section IV. Our predictions for the partial decay widths of these modes are collected in Table 1. The width of the molecule $\mathcal{M}_{\mathrm{AV}}$ in this case is $\Gamma\left[\mathcal{M}_{\mathrm{AV}}\right]=(30\pm 4)~\mathrm{MeV}$, which is not small and determines the mean lifetime of this structure.
[…]
$B^{\ast-}\overline{D}^{0}$ $13.2\pm 2.8$ $7.9\pm 2.4$
$B^{-}\overline{D}^{\ast 0}$ $3.76\pm 0.75$ $3.9\pm 1.1$
$\overline{B}^{\ast 0}_{s}D^{-}_{s}$ $9.34\pm 1.75$ $3.8\pm 1.0$
$\overline{B}_{s}^{0}D_{s}^{\ast-}$ $3.07\pm 0.62$ $2.3\pm 0.7$
Table 1: Decay modes of the molecule $\mathcal{M}_{\mathrm{AV}}$ at the lower limit of its mass $15710~\mathrm{MeV}$, relevant strong couplings, and partial widths $\Gamma$.
Investigations performed in the present work provide interesting information on parameters of the fully heavy axial-vector hadronic molecules, but, at the same time, give rise to different questions which should be clarified in the context of presumably alternative models and methods.